OpenStudy (anonymous):

What are the exact solutions of x2 - 5x - 1 = 0? Answer x = the quantity of negative 5 plus or minus the square root of 21 all over 2 x = the quantity of 5 plus or minus the square root of 29 all over 2 x = the quantity of 5 plus or minus the square root of 21 all over 2 x = the quantity of negative 5 plus or minus the square root of 29 all over 2

OpenStudy (mimi_x3):

Use the quadratic formula: \[\large x=\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} \] & sub it it in: \[a = 1 ; b =-5;c=-1\]

hero (hero):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Mine looks better :P

OpenStudy (mimi_x3):

\[\color{ red} {x=\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} }\] my one has colour now :P

OpenStudy (anonymous):